### v2

# Rethinking Pseudo-Label Guided Learning for Weakly Supervised Temporal Action Localization from the Perspective of Noise Correction

Abstract
Pseudo-label learning methods have been widely applied in weakly-supervised temporal action localization. Existing works directly utilize weakly-supervised base model to generate instance-level pseudo-labels for training the fully-supervised detection head. We argue that the noise in pseudo-labels would interfere with the learning of fully-supervised detection head, leading to significant performance leakage. Issues with noisy labels include:(1) inaccurate boundary localization; (2) undetected short action clips; (3) multiple adjacent segments incorrectly detected as one segment. To target these issues, we introduce a two-stage noisy label learning strategy to harness every potential useful signal in noisy labels. First, we propose a frame-level pseudo-label generation model with a context-aware denoising algorithm to refine the boundaries. Second, we introduce an online-revised teacher-student framework with a missing instance compensation module and an ambiguous instance correction module to solve the short-action-missing and many-to-one problems. Besides, we apply a high-quality pseudo-label mining loss in our online-revised teacher-student framework to add different weights to the noisy labels to train more effectively. Our model outperforms the previous state-of-the-art method in detection accuracy and inference speed greatly upon the THUMOS14 and ActivityNet v1.2 benchmarks.
Introduction
Temporal action localization (TAL) (Chao et al. 2018) is a significant task in video comprehension, which aims to locate the start and end times of target action instances within untrimmed videos and identify their categories. It has shown great potential in tasks like intelligent surveillance, video summarization, etc. Inspired by the widespread application of deep learning methods across various fields(Chen et al. 2024; Zhang et al. 2024b; Lu et al. 2024b, a, c), temporal action localization has widely adopted deep learning-based approaches, which has sparked significant attention. Although the TAL problem has garnered much attention, current prominent methods often rely on costly, labor-intensive temporal boundary annotations, hindering the widespread applications of TAL.
To reduce reliance on instance-level labels, Weakly Supervised Temporal Action Localization (WTAL) solely based on video-level labels (Chen et al. 2022; Zhang and Qi 2024) has gained increasing attention. WTAL only requires knowledge of which actions occurred in the video, without the need for precise location information.
Figure 1: Noise correction modeling for WTAL. NoCo introduces pseudo-label noise correction modules to address the typical failures in existing WTAL methods.
Existing WTAL works mainly follow per-category localization pipelines (Wang et al. 2017; Zhao et al. 2022), using video-level classifiers trained on class labels (Narayan et al. 2021) to filter each video clip. However, due to the absence of refinement mechanisms, these models attribute high confidence to incorrect segments, such as background elements, leading to imprecise localization. Many studies (Liu, Jiang, and Wang 2019; Liu et al. 2021a, b) attempt to address the disparity between classification and localization, with a promising solution being the generation and utilization of pseudo-labels which benefits from utilizing clip-level supervision instead of video-level labels. Existing works (Yang et al. 2021c; Luo et al. 2020; Pardo et al. 2021; Zhai et al. 2020) have achieved remarkable results by introducing pseudo-labels on WTAL.
However, due to the sparsity of video-level label supervision information, there exists a large amount of noises in pseudo-labels. We argue that these noises can be classified into three types: 1) Inaccurate instance boundaries; 2) Missing action instances; 3) Adjacent instances aggregated into one instance, as shown in Figure 1.
To mitigate the impact of noisy pseudo-labels on WTAL, we reevaluate the pseudo-label learning process from the perspective of noise correction. We propose NoCo, a weakly-supervised temporal action localization method, which trains a weakly-supervised pseudo-label generator first, and then incorporate proposed teacher-student-based online noise learning framework over noisy pseudo-labels. In practice, there exists two issues in the above two stages. (1) How to generate accurate pseudo-label. (2) How to implement the online noise correction framework.
To generate accurate pseudo-labels, we argue that using Non-Maximum Suppression (NMS) (Neubeck and Van Gool 2006) loses the contextual information, leading to inaccurate instance boundaries. In the pseudo-label generation process, contextual information serves as a valuable supplementary source. Motivated by this, we introduce the Context-Aware Label Augmentation module (CALA), which adaptively perceives the contextual information for each pseudo-label and and corrects them through weighted averaging. Such augmentation not only yields more accurate action boundaries but also provides a more reasonable initial pseudo-label set for subsequent noise correction framework.
To realize online noise correction, we introduce a strategy that combines noise correction with teacher-student training. This approach leverages the potential of the teacher-student framework in noise learning, allows for the convenient addition of modules tailored for WTAL, and ensures high inference efficiency. Different from the previous teacher-student strategy, our approach introduces several key innovations. Firstly, to generate robust and stable pseudo-label compensation information, our teacher model employs weighted aggregation of multiple historical student models. Secondly, we propose a Missing Instance Compensation module (MIC) to solve the short clips missing problem by identifying compensation instances with low intersection-over-union (IoU) ratio from the teacher prediction information. Thirdly, the Ambiguous Instance Correction module (AIC) is introduced to address the issue of aggregating adjacent instances into one by mining the high IoU ratio instance from the teacher’s online pseudo-labels. Lastly, a High-quality Pseudo-label Mining loss (HPM) is proposed to construct adaptive optimization weights, fully leveraging high-quality supervision information.
It should be noted that our noise correction framework is decoupled from the WTAL base model. The inference process of NoCo does not depend on the WTAL base model.In this paper, we build a student model based on the efficient TAL model TriDet (Shi et al. 2023), and during the inference phase, only the student model needs to be run.
In summary ,the main contributions are listed as follows.
  • We revisit the WTAL process from the perspective of noise correction and propose NoCo, a noise learning framework aimed at gradually mitigating the pseudo-label noise effect in TAL. We observe three types of noise information present in the pseudo-label learning method.
  • We propose a Context-aware Label Augmentation module aimed at rectifying inaccurate action boundaries. We present a teacher-student-based online noise correction framework to address short-action omissions and many-to-one problems, encompassing Missing Instance Compensation, Ambiguous Instance Correction , and High-quality Pseudo-label Mining.
  • We conduct extensive experiments to demonstrate the effectiveness of NoCo. Our methods bring a significant improvement for WTAL baselines on two benchmarks and achieve state-of-the-art performance. Additionally, NoCo demonstrates fast inference speed.
Related Work
Weakly-supervised temporal action localization aims to determine the start and end times, as well as the category, of specified action instances in untrimmed videos using only video-level labels. Unlike traditional fully-supervised methods, it doesn’t rely on costly instance-level labels, making it a widely studied approach(Paul, Roy, and Roy-Chowdhury 2018). Existing weakly-supervised methods fall into four categories: a) Multi-instance Learning: UntrimmedNet(Wang et al. 2017) was first applied to WTAL, generating action proposals which were then classified. However, such methods tend to focus on the most discriminative regions due to the difference between classification and localization tasks. b) Metric Learning: WTALC(Paul, Roy, and Roy-Chowdhury 2018) utilizes metric learning to make features of the same class similar. c) Attention-Based Learning: STPN(Nguyen et al. 2018) introduced a sparse loss for attention sequences, capturing key foreground segments. d) Pseudo-label Learning: TSCN(Zhai et al. 2020) generates segment-level pseudo-labels by fusing attention from RGB and FLOW modalities,while UGCT (Yang et al. 2021c) considers their complementarity and ASM-Loc (He et al. 2022) enhances features with action proposals. Previous pseudo-label learning methods have made great progress, but did not consider the existence of noise in pseudo-labels.We are the first to improve pseudo-label learning methods from the perspective of noise correction.
Figure 2: (a). Framework Overview. The dark gray modules indicate the base method (eg. WTAL model and augmentation), while the others are our noise correction modules. (b). Ambiguous Instances Correction is based on teacher predictions to accurate instance boundaries through mining high IoU samples and adaptively aggregating context. It is able to solve many-to-one and in-accurate boundary position problem. (c). Missing Instance Compensation focuses on adding missing instance based on teacher predictions. It aims to solve short action missing problem. (d). High-Quality Pseudo-label Mining Loss assigns adaptive weights for each noisy action instance to mine high-quality pseudo-label.
Pseudo Label Guided Training has been widely applied in visual tasks with weak or limited supervision. In weakly-supervised object detection, one common approach is self-training (Ren et al. 2020). This method involves training an teacher model first and using high-confidence predictions to generate instance-level pseudo-labels. These pseudo-labels are then utilized to train the final detector. Similarly, in semi-supervised learning (Weng et al. 2022) and domain adaptation (Das and Lee 2018), models are first trained on labeled/source datasets and then used to generate pseudo-labels for unlabeled/target datasets, guiding the training process. Similar to these works, our method utilizes segment-level pseudo-labels to guide training process for the WTAL task. The difference is that we notice noises in the pseudo-label, and progressively correct these noises.
Learning with Noisy Labels has been widely applied in image classification tasks. To effectively leverage noise labels to guide the training process, DivideMix (Li, Socher, and Hoi 2020) uses two networks for sample selection via binary mixture modeling. Price et al. (Pleiss et al. 2020) introduced the marginal area statistic, measuring the average difference between a sample’s assigned class logit and its highest unspecified class logit, to separate correctly from incorrectly labeled data. Liu et al. (Liu et al. 2020) found the model initially predicts ground-truth labels but eventually memorizes wrong ones, motivating exploiting the model’s early output. Motivated by these works, we propose noisy label learning framework for TAL task.
Methods
In this section, we introduce the proposed NoCo framework in detail. First, we present the problem definition and describe the framework overview. Then we present the main components of NoCo in detail.
Problem Definition. Given an untrimmed training video $V_{i}$ and its video-level label $y_{i}\in\{0,1\}^{C}$, where $y_{i}$ is an C-dimension one-hot vector indicating category. The goal of the inference phase is to predict an action instance set $\mathcal{A}=\{(c_{i},\theta_{i},s_{i},e_{i})\}$, where $c_{i}$ denotes the category of the instance $A_{i}\in\mathcal{A}$, $\theta_{i}$ is the confidence score, and $s_{i}$ and $e_{i}$ denote the start and end timestamp.
Framework Overview. As illustrated in Figure 2, our noise correction framework comprises two main stages. Firstly, the WTAL Base Model generates noisy pseudo-labels and then CALA is employed for context-aware enhancement of noisy pseudo-labels to achieve more accurate boundaries. Secondly, the enhanced pseudo-labels are input into the online noise correction framework as the initial pseudo-label set.
In the online correction framework, the teacher model is aggregated from various historical versions of student models, providing stable compensatory information for noise correction. The Missing Instance Compensation module (MIC) identifies action instances with low IoU with the pseudo-label, generating compensatory action instances. Simultaneously, the Ambiguous Instance Correction module (AIC) discovers action instances with high IoU with the pseudo-label, creating an associated action instance set. Then this associated set is used to rectify over-complete action instances in pseudo-labels, aligning them with corrected instances. Adjacent missing instances are compensated through MIC. The High-Quality Pseudo-Label Mining Loss (HPM), combined with the associated instance set, adaptively mines high-quality pseudo-labels, assigning greater optimization weights to these labels.
Noisy Label Generator
We first train the WTAL model based on video-level labels to obtain instance-level pseudo-labels. To refine the boundary of inaccurate instance of pseudo-labels, we propose CALA to enhance localization through context augmentation. The detail is shown as follows.
WTAL Base Model
Following the previous work (Pardo et al. 2021; He et al. 2022), for each video $V_{i}$, we use the features $F\in R^{2048}$ extracted by the I3D network (Carreira and Zisserman 2017), where the I3D network is pre-trained on the Kinetics400 dataset. Using video-level labels, we train a WTAL via which we can obtain a noisy pseudo-label for each video sample. In this paper, the WTAL we use is ASM-Loc (He et al. 2022) if not specified. Note that our method is decoupled from the WTAL model and can be easily cooperated to other weakly-supervised models. In the experiment section, we conduct an analysis of the generalizability for WTAL models.
Context-Aware Label Augmentation
Due to the sparsity of video-level label information, there exists non-negligible noises in the obtained pseudo-labels.
We propose the Context-Aware Label Augmentation (CALA) algorithm to refine instance pseudo-labels and reduce noise by considering neighboring instances. The CALA process is outlined in Algorithm 1.
Algorithm 1 Context-Aware Label Augmentation (CALA)
Input: Weakly-supervised output $\mathcal{P}_{w}$, NMS threshold $\rho$, Confidence threshold $\psi$, IoU threshold $\eta_{0}$, Weight Aggregation Function $\mathcal{F}_{fuse}^{aug}(\cdot$, $\cdot$)⋅ )
1:  $\mathcal{A}^{aug},\mathcal{W}^{aug}=\phi$
2:  $\mathcal{P}_{filt}^{aug}=\{p_{t}|p_{t}\in\mathcal{P}_{w}\;\&\;{p_{t}}(\theta_{p_{t}})>\psi\}$
3:  $index_{keep}^{aug}$=NMS$(\mathcal{P}_{filt}^{aug},\rho)$
4:  $\mathcal{P}_{keep}^{aug}=\mathcal{P}_{filt}^{aug}[index_{keep}^{aug},:]$
5:  for every $P_{a}\in\mathcal{P}_{keep}^{aug}$ do
6:     $\mathcal{P}_{a,ass},\mathcal{W}_{a,ass}=\phi$
7:     for every $P_{w}\in\mathcal{P}_{w}$ do
8:        if $P_{a}(c_{a})=P_{w}(c_{w})$ and IoU($P_{a},P_{w}$)$>\eta_{0}$ then
9:           # every associated action instance of $P_{a}$
10:           Calculate adaptive weights $w_{a,w}$
11:           $\mathcal{P}_{a,ass}=\mathcal{P}_{a,ass}\cup P_{w}$, $\mathcal{W}_{a,ass}=\mathcal{W}_{a,ass}\cup w_{a,w}$
12:        end if
13:     end for
14:     if $\mathcal{P}_{a,ass}\text{is not}\;\phi$ then
15:        $P_{a}^{aug},w_{a}^{aug}=\mathcal{F}_{fuse}^{aug}(\mathcal{P}_{a,ass},\mathcal{W}_{a,ass}$)
16:     end if
17:     $\mathcal{A}^{aug}=\mathcal{A}^{aug}\cup P_{a}^{aug}$, $\mathcal{W}^{aug}=\mathcal{W}^{aug}\cup w_{a}^{aug}$
18:  end for
19:  return  Augmented Pseudo Ground-Truth $\mathcal{A}^{aug}$ and Weight $\mathcal{W}^{aug}$.
First, initialize the enhanced pseudo-labels and their corresponding weights. By setting a confidence threshold and deploying NMS processing, a set of high-confidence and low-overlapping action instances $P_{keep}^{aug}$ is selected. The context enhancement process is executed for each action instance $P_{a}$ in $\mathcal{P}_{keep}^{aug}$, including determining the associated action instance set $\mathcal{P}_{a,ass}$ and calculating the adaptive weights. The constraint for the associated action instance is: a) sharing the same category; b) having an IoU greater than a given threshold $\eta_{0}$. The formula for calculating the adaptive weight of $P_{a}$’s associate instance $P_{r}$ is as follows:
$w_{r}=e^{\sqrt{\text{IoU}(P_{r},\mathcal{P}_{keep}^{aug})\cdot\text{min}(\text{max}(P_{r}(\theta_{r}),0),1)}},$ (1)
where $P_{r}(\theta_{r})$ is the confidence score of $P_{r}\in\mathcal{P}_{a,ass}$. If $\mathcal{P}_{a,ass}$ is not empty, then the augmented $P_{a}^{aug}$ of $P_{a}$ is obtained via weighted aggregation of associated action instances set $\mathcal{P}_{a,ass}$, which is computed as follows:
$P_{a}^{aug}(\zeta_{a})$ $=\mathcal{F}_{fuse,stage}^{aug}(P_{a},\mathcal{P}_{a,ass})$ (2)
$=\frac{\sum_{i=1}^{n}w_{i}\cdot P_{i}(\zeta_{i}),P_{i}\in\mathcal{P}_{a,ass}}{\sum_{i=1}^{n}w_{i}},$
$\text{where}\;\zeta_{i}=\left\{\begin{array}[]{rcl}s_{i}&&stage=start\\ e_{i}&&stage=end,\\ \end{array}\right.$ (3)
weight aggregation function $\mathcal{F}_{fuse}^{aug}(\cdot$, $\cdot$)⋅ ) consists of two stage: start timestamp aggregation $\mathcal{F}_{fuse,start}^{aug}(\cdot$, $\cdot$)⋅ ) and end timestamp aggregation $\mathcal{F}_{fuse,end}^{aug}(\cdot$, $\cdot$)⋅ ). The Weight of $P_{a}^{aug}$ is computed as $w_{a}^{aug}=\sqrt[n]{\prod_{i=1}^{n}w_{i}}.$ Then add the $P_{a}^{aug}$ to the augmented pseudo-label set $\mathcal{A}^{aug}$ and repeat this process till all action instance is scanned.
Online Noise Correction Framework
The online noise correction framework aims to progressively correct three types of noise: missing action instances, many-to-one and incomplete instance problem. Specifically, we introduce a teacher-student-based online noise correction framework. As illustrated in Figure 2, our online noise correction framework consists of several components: feature extraction encoder, data augmentation module, student model, teacher model, AIC, MIC, online pseudo-label set $\mathcal{A}^{*}$ and HPM.
Teacher Model Pre-training
The online noise correction process begins with the pre-training of the teacher model, involving using the enhanced pseudo-labels generated by CALA for full supervised training. After completing pre-training, the teacher model plays a crucial role in the online generation of stable compensation information during the noise correction stage, aiding the processes of AIC and MIC. The online pseudo-label $\mathcal{A}^{*}$ and weight $\mathcal{W}^{*}$ is first initialized as $\mathcal{A}^{aug}$ and $\mathcal{W}^{aug}$.
Ambiguous Instances Correction
Ambiguous Instances Correction (AIC) aims to leverage the pseudo-label information derived from teacher predictions to accurately position instance boundaries through adjacent context aggregation, thus enabling the online update of the pseudo-label set. For low-quality samples, the output from the teacher model is further corrected through the AIC module, effectively suppressing noise in the pseudo-labels. Compared with traditional semi-supervised learning, we possess not only the output of the teacher model but also maintain an online, real-time corrected pseudo-label set. We refer to Supplementary for the workflow of AIC.
The input to the AIC consists of teacher predictions $\mathcal{P}_{w}$ and the online pseudo-label set $\mathcal{A}^{*}$. For each instance $A_{r}^{*}$ in the $\mathcal{A}^{*}$, we calculate its IoU with every action instance $P_{t}\in\mathcal{P}_{w}$. If the IoU is greater than the threshold $\eta_{1}$, and the confidence score of the action instance exceeds the threshold $\psi$, we include it in the corresponding associated set $\mathcal{A}_{r,ass}$ and compute its adaptive weight. The formula for calculating the adaptive weight $w_{t}$ of $A_{r}^{*}$’s associated instance $P_{t}$ is the same as the one proposed in the CALA algorithm, as described above in equation (1). By combining $\mathcal{A}_{r,ass}$ and $A_{r}^{*}$, we perform weighted aggregation to obtain the corrected action instance $A_{r}^{cor}$ of $A_{r}^{*}$ as follows:
$A_{r}^{cor}(\zeta_{r})$ $=\mathcal{F}_{fuse,stage}^{cor}(A_{r}^{*},\mathcal{A}_{r,ass})=\alpha\cdot w_{r}^{*}\cdot A_{r}^{*}(\zeta_{r})$ (4)
$+(1-\alpha)\cdot\frac{\sum_{i=1}^{n}w_{i}\cdot A_{i}(\zeta_{i}),A_{i}\in\mathcal{A}_{r,ass}}{\sum_{i=1}^{n}w_{i}},$
$\text{where}\;\zeta_{i}=\left\{\begin{array}[]{rcl}s_{i}&&stage=start\\ e_{i}&&stage=end,\\ \end{array}\right.$ (5)
where $\alpha$ is hyper-parameter, $w_{r}^{*}$ is the weight of instance $A_{r}^{*}$’s weight in $\mathcal{W}^{*}$, $n$ is the size of $A_{r,ass}$.The corrected weight of $A_{r}^{*}$ is computed as:
$w_{r}^{cor}=\sqrt[n+\beta]{(\prod_{i=1}^{n}w_{i})\cdot(w_{r}^{*})^{\beta}},\;\;A_{i}\in\mathcal{A}_{r,ass},$ (6)
where $\beta$ is hyper-parameter. Then add the corrected instance $A_{r}^{aug}$ of $A_{r}^{*}$ to the corrected pseudo-label set $\mathcal{A}^{cor}$, and add $w_{r}^{cor}$ to corrected weight set $\mathcal{W}^{cor}$.
When all instance in $\mathcal{A}^{*}$ is corrected, the online set $\mathcal{A}^{*}$ is updated as $\mathcal{A}^{cor}$. The online weight $\mathcal{W}^{*}$ is updated as $\mathcal{W}^{cor}$. Through this updating, the $\mathcal{A}^{*}$ and $\mathcal{W}^{*}$ is corrected.
Missing Instance Compensation
MIC aims to further update the pseudo-label set online, leveraging the potential missing information in neighboring pseudo-label instances, which are generated based on teacher predictions. Although AIC effectively utilizes contextual information within the pseudo-labels, it only focuses on correcting existing action instances in noisy pseudo-labels and does not address the task of compensating missing action instances. We refer to Supplementary for the main workflow of the MIC.
MIC takes teacher predictions $\mathcal{P}_{t}$ and the online pseudo-label set $\mathcal{A}^{*}$ as inputs. First, MIC filters out action instances with low confidence from $\mathcal{P}_{t}$, saving as $\mathcal{P}_{filt}^{com}$. Then it applies NMS to select low IoU instance on $\mathcal{P}_{filt}^{com}$ to obtain $\mathcal{P}_{keep}^{com}$. Next, for each instance $P_{n}$ in $\mathcal{P}_{keep}^{com}$, MIC computes the IoU with every action instance in $\mathcal{A}^{*}$. If the maximum IoU is less than $\eta_{2}$, $P_{n}$ is included in the compensated pseudo-label set $\mathcal{A}^{com}$. The adaptive weight $w_{n}^{com}$ of $A_{n}\in\mathcal{A}^{com}$ is obtained through:
$w_{n}^{com}=e^{\text{min}(\text{max}(A_{n}(\theta_{n}),0),1)},$ (7)
where $A_{n}(\theta_{n})$ is the confidence score of $A_{n}$. $w_{n}^{com}$ is added into compensated weight $\mathcal{W}^{com}$.
Through the MIC module, missing action instances are filled. Similar before, we update $A^{*}$ as $\mathcal{A}^{com}$, $\mathcal{W}^{*}$ as $\mathcal{W}^{com}$ again to achieve online label compensation.
High-Quality Pseudo-Label Mining
To fully utilize the supervision information from accurate pseudo-labels, we introduce adaptive weights for each action instance by assigning higher optimization weights to high-quality instances and vice versa. Specially, we propose the High-Quality Pseudo-Label Mining (HPM) Loss, which is formulated as:
$L$ $=\frac{\lambda}{N_{pos}}\sum_{l,t}\mathds{1}_{\{c_{t}^{l}>0\}}\cdot w_{t}\cdot(\sigma_{IoU}L_{cls}+L_{reg})$ (8)
$+\frac{1}{N_{reg}}\sum_{l,t}\mathds{1}_{\{c_{t}^{l}=0\}}\cdot L_{cls},$
where $\lambda$ represents the weight of positive sample loss, $N_{pos}$ and $N_{neg}$ denote the number of positive and negative samples, and $l$ stands for the number of classification heads and Trident-head heads. For each classification head and Trident-head, we compute foreground classification loss $L_{cls}$, regression loss $L_{reg}$, and background classification loss sequentially. In this paper, $L_{cls}$ is focal loss and $L_{reg}$ is IoU loss. It’s worth noting that using adaptive weights $w_{t}$ for each foreground action instance gives more importance to high-quality instances during the optimization process.
mAP@IoU(%) AVG mAP(%)
Sup. Method 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.1:0.5 0.3:0.7 0.1:0.7
SSN (Zhao et al. 2017) 60.3 56.2 50.6 40.8 29.1 - - 47.4 - -
Fully TAL-Net(Chao et al. 2018) 59.8 57.1 53.2 48.5 42.8 33.8 20.8 52.3 39.8 45.1
SF-Net (Ma et al. 2020) 68.3 62.3 52.8 42.2 30.5 20.6 12.0 51.2 31.6 41.2
DCM (Ju et al. 2021) 70.2 63.5 55.6 44.7 32.3 22.0 12.3 53.3 33.4 42.9
BackTAL(Yang et al. 2021a) - - 54.4 45.5 36.3 26.2 14.8 - 35.4 -
PCL (Li, Cao, and Ye 2023) 74.6 70.2 63.3 55.9 44.4 - - 61.7 - -
LACP (Lee and Byun 2021) 75.7 71.4 64.6 56.5 45.3 34.5 21.8 62.7 44.5 52.8
LACP+NoCo 78.8 74.6 68.9 60.1 49.6 37.2 24.2 66.4 48.0 56.2
HR-Pro (Zhang et al. 2024a) 85.6 81.6 74.3 64.3 52.2 39.8 24.8 71.6 51.1 60.4
Point-level HR-Pro+NoCo 86.0 82.1 75.6 65.7 53.6 40.6 25.3 72.6 52.2 61.3
CoLA (Zhang et al. 2021) 66.2 59.5 51.5 41.9 32.2 22.0 13.1 50.3 32.1 40.9
DELU (Chen et al. 2022) 71.5 66.2 56.5 47.7 40.5 27.2 15.3 56.5 37.4 46.4
Zhou et al. (Zhou et al. 2023) 74.0 69.4 60.7 51.8 42.7 26.2 13.1 59.7 38.9 48.3
Ju et al. (Ju et al. 2023) 73.5 68.8 61.5 53.8 42.0 29.4 16.8 59.9 40.7 49.4
PivoTAL (Rizve et al. 2023) 74.1 69.6 61.7 52.1 42.8 30.6 16.7 60.1 40.8 49.7
CASE (Liu et al. 2023) 72.3 67.1 59.2 49.4 37.7 24.2 13.7 57.1 36.8 46.2
ISSF (Yun et al. 2024) 72.4 66.9 58.4 49.7 41.8 25.5 12.8 57.8 37.6 46.8
CO2-Net (Hong et al. 2021) 70.1 63.6 54.5 45.7 38.3 26.4 13.4 54.4 35.7 44.6
CO2-Net+NoCo 74.6 70.3 63.0 53.9 44.0 31.7 17.4 61.2 42.0 50.7
ASM-Loc (He et al. 2022) 71.2 65.5 57.1 46.8 36.6 25.2 13.4 55.4 35.8 45.1
Video-level ASM-Loc+NoCo 75.2 70.7 63.0 54.1 44.0 31.7 17.7 61.4 42.1 50.9
Table 1: Comparisons with state-of-the-art methods. AVG is the average mAP under the IoU thresholds 0.1:0.7:0.1 for THUMOS14 dataset. ’-’ means that the corresponding results are nor reported in the original papers.
mAP@IoU(%)
Sup. Method 0.5 0.75 0.95 AVG
Fully SSN (Zhao et al. 2017) 41.3 27.0 6.1 26.6
SF-Net (Ma et al. 2020) 37.8 - - 22.8
BackTAL (Yang et al. 2021a) 41.5 27.3 4.7 27.0
BackTAL+NoCo 49.8 31.6 6.1 30.8
LACP (Lee and Byun 2021) 44.0 26.0 5.9 26.8
Point-level LACP+NoCo 50.2 30.1 6.8 30.7
CoLA (Zhang et al. 2021) 42.7 25.7 5.8 26.1
DELU (Chen et al. 2022) 44.2 26.7 5.4 26.9
Ren et al. (Ren et al. 2023) 44.2 26.1 5.3 25.5
Ju et al. (Ju et al. 2023) 48.3 29.3 6.1 29.6
CASE (Liu et al. 2023) 43.8 27.2 6.7 27.9
CO2-Net (Hong et al. 2021) 43.3 26.3 5.2 26.4
CO2-Net+NoCo 48.9 30.0 6.6 30.1
ASM-Loc (He et al. 2022) 43.4 26.6 5.3 26.5
Video-level ASM-Loc+NoCo 49.1 30.5 6.7 30.7
Table 2: Detection results on ActivityNet1.2
# CALA T-S AIC MIC HPM AVG mAP(%)
a 35.8
b $\checkmark$ 37.1 (+1.3)
c $\checkmark$ $\checkmark$ 37.8 (+2.0)
d $\checkmark$ $\checkmark$ $\checkmark$ 40.0 (+4.2)
e $\checkmark$ $\checkmark$ $\checkmark$ 39.6 (+3.8)
f $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 41.9 (+6.1)
g $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 42.1 (+6.3)
Table 3: Ablation study of core components,’T-S’ denotes the Teacher-Student module.
WTAL CALA T-S MIC AIC HPM AVG mAP(%)
UM 30.3
$\checkmark$ 31.6 (+1.3)
$\checkmark$ $\checkmark$ 32.3 (+2.0)
$\checkmark$ $\checkmark$ $\checkmark$ 33.9 (+3.6)
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 36.1 (+5.8)
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 36.4 (+6.1)
CO2-Net 36.0
$\checkmark$ 37.0 (+1.0)
$\checkmark$ $\checkmark$ 37.7 (+1.7)
$\checkmark$ $\checkmark$ $\checkmark$ 40.2 (+4.2)
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 41.8 (+5.8)
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 42.0 (+6.0)
Table 4: Performances under two typical WTAL algorithms on THUMOS14 dataset.
Experiment
Settings and Datasets
Datasets. We evaluate our method on THUMOS14 and ActivityNet v1.2, following (Yang et al. 2021b). THUMOS14 contains 200 validation and 213 test samples. The validation set is used for training, and the test set for evaluation. ActivityNet v1.2 includes 4,819 training, 2,383 validation, and 2,489 test videos across 100 action classes, averaging 1.5 instances per video. We use the training set for training and the validation set for evaluation.
Implementation.We use the I3D model for feature extraction and ASM-Loc (He et al. 2022) as the default noise label generator unless stated otherwise. Additional dataset-specific details are in the supplementary material. We follow standard protocols with IoU thresholds to measure localization accuracy and report mIoU for pseudo-label quality across foreground and background categories.
Comparison with State-Of-Art Methods
We compare NoCo with the existing methods on ActivityNet v1.2 and THUMOS14.The result is shown in Table 1.
Results on THUMOS14. NoCo outperforms all WTAL baseline methods. Its average mAP is 6.1% higher than CO2-Net and 5.8% higher than ASM-Loc . Additionally, NoCo surpasses the latest methods CASE , AHIM , DDG-Net , and the current leading method PivoTAL, pushing the average mAP to 50.9%, setting a new state-of-the-art performance. Even when compared with FTAL methods, NoCo outperforms TAL-Net at all IoUs. Furthermore, we also applied NoCo to point-supervised temporal action localization methods LACP and HR-Pro, achieving 3.4% and 0.9% average mAP improvements over the baselines, respectively, resulting in current state-of-the-art performance.
Results on ActivityNet v1.2. Our method outperforms existing methods across all metrics. Compared to the baseline ASM-Loc, our method improves mAP by 4.2%, and compared to the baseline CO2-Net, it improves by 3.7%. NoCo also surpasses Ju et al.(Ju et al. 2023), achieving a new state-of-the-art localization accuracy of 30.7%. Our method even outperforms some previous fully supervised methods, such as SSN, demonstrating the feasibility of noise correction modeling. Additionally, we applied NoCo to point-supervised temporal action localization methods BackTAL and LACP, achieving average mAP improvements of 3.8% and 3.9%, respectively, compared to the baseline.
The consistent results on ActivityNet v1.2 and THUMOS14 demonstrate the effectiveness of our NoCo.
Ablation Studies
We conducted a series of ablation experiments to evaluate the contributions of each component in our method. The results, presented in Table 3, are labeled from a to g for clarity.
Baseline. Experiment a combines the weakly supervised model ASM-Loc with the fully supervised TriDet model for pseudo-label retraining on the THUMOS14 dataset, without any correction modules. Pseudo-labels were generated by filtering action instances with confidence scores above 0.5. However, this threshold-based method introduced substantial noise, resulting in minimal performance improvement.
Figure 3: Visualization of ground-truth and predictions.
Influence of CALA. Experiment b applies CALA, achieving a 1.3% mAP increase, demonstrating its ability to enhance pseudo-labels with contextual information.
Influence of Customized Teacher-Student Training. Experiment c introduces a customized teacher-student framework without pseudo-label correction modules. The teacher’s labels guide the student, improving performance by 0.7% due to the student model’s robustness from prior teacher-student iterations.
Influence of AIC. Experiment d adds AIC, boosting performance by 2.2%. AIC corrects inaccurately localized pseudo-labels, refining the student model’s learning.
Influence of MIC. Experiment e incorporates MIC, resulting in a 1.8% improvement. MIC compensates for missing action instances in pseudo-labels, enhancing learning.
Influence of Combined MIC and AIC. Experiment f applies both MIC and AIC, yielding a 4.1% performance increase over experiment c. AIC corrects overly complete instances, while MIC fills in missing actions, addressing the one-to-many issue.
Influence of High-Quality Pseudo-Label Mining. Experiment g adds high-quality pseudo-label mining loss, providing a 0.2% performance boost and reaching the highest mAP of 42.1%.
Generalizability Analysis of WTAL Models
Our noise correction framework is decoupled from weakly supervised models, making it applicable to any existing temporal action localization method. To evaluate its generalizability, we apply our framework to the UM (Lee et al. 2021) and CO2-Net (Hong et al. 2021) models, with results shown in Table 4.
For the UM model, the average mAP on THUMOS14 is 30.3%. The CALA module improves this by 1.3%, and the Teacher-Student module contributes an additional 2.0%. The MIC, AIC, and HPM modules further increase the performance by 3.6%, 5.8%, and 6.1%, respectively.
Similarly, for CO2-Net, the mAP starts at 36.0%. The sequential introduction of the CALA, Teacher-Student, MIC, AIC, and HPM modules results in improvements of 1.0%, 1.7%, 4.2%, 5.8%, and 6.0%, respectively.
These results highlight the strong generalizability of our noise correction framework across various weakly-supervised models.
Analysis of Inference Speed
NoCo gains a speed advantage by running only the student model during inference. We compare the inference performance and speed in Table 5. It can be observed that NoCo is not only faster than the baseline WTAL method, showing a threefold improvement, but also achieves more accurate localization results.
mAP(0.3:0.7)↑ speed(video/s)↑ FLOPS↓
Baseline 35.8 1.64 124.79G
NoCo 42.1 6.67 20.12G
Table 5: NoCo outperforms existing methods in terms of both localization average mAP and inference speed.
Qualitative Results
Figure 3 illustrates the visual comparison between the base model and NoCo for the action Playing congas. We observe that when introducing AIC, the action boundary is more precise compared to the baseline. Furthermore, with the addition of MIC, missing action instances are compensated for, and many-to-one problem is solved. These results provide evidence for the effectiveness of proposed AIC and MIC.
Conslusion
In this paper, we propose a progressive framework for correcting noisy pseudo-labels. We introduce a CALA module to obtain more accurate action boundaries. Subsequently, we establish an online noise correction framework based on a teacher-student training strategy. This framework incorporates the AIC and MIC module to tackle the instances-missing and many-to-one problem. Furthermore, we introduce an HPM loss to discover high-quality pseudo-labels. Our approach achieves state-of-the-art performance on THUMOS14 and ActivityNet v1.2, with a threefold improvement in inference speed.
Acknowledgments
This work was supported by the National Key R&D Program of China (2022YFB4701400/4701402), SSTIC Grant (KJZD20230923115106012,KJZD20230923114916032,
GJHZ20240218113604008), Beijing Key Lab of Networked Multimedia and National Natural Science Foundation of China under Grant 62202302.